# Stolz positive scalar curvature structure groups, proper actions and equivariant 2-types

### v2

@@ 3 The Stolz exact sequence
  1. 1.
    For every component $X_{c}$ of $X$ consider the induced $G$-invariant subspace $G\cdot X_{c}$. We then have a disjoint union decomposition into $G$-invariant subsets $X=\bigsqcup G\cdot X_{c}$ with $X_{c}$ connected (choosing one representative in each $G$-orbit of components).
    Clearly, every group in the Stolz exact sequence and the whole sequence then split canonically as a direct product, with one factor for each $G\cdot X_{c}$.
  2. 2.
    For a $G$-space $Y=G\cdot X_{c}$ with $X_{c}$ connected, consider the subgroup $G_{0}\subset G$ of all elements which map $X_{c}$ to itself. Then $X_{c}$ is a $G_{0}$-space and $Y$ is obtained by “induction”: $Y=G\times_{G_{0}}X_{c}$. Whenever we have a $G$-map $f\colon M\to Y$ which could be part of a cycle or a bordism for the groups in the Stolz exact sequence, then we obtain $M_{0}:=f^{-1}(X_{c})$ a union of components of $M$ on which $G_{0}$ acts (restricting the action of $G$ on $M$). Then $f|_{M_{0}}\colon M_{0}\to X_{c}$ is a $G_{0}$-equivariant map and we obtain $M$ and $f$ by induction: $M=G\times_{G_{0}}M_{0}$ and $f=id_{G}\times_{G_{0}}f|_{M_{0}}$.
    It follows that induction gives an isomorphism (already on the level of cycles and relations)
    $\mathrm{ind}_{G_{0}}^{G}\colon R^{spin}_{*}(X_{c})^{G_{0}}\to R^{spin}_{*}(Y)^{G}$
    and the same for the other groups in the sequence 3.3 and the maps between them.
@@
This again follows from covering theory: if we fix $x\in X$ with lift $\tilde{x}\in\widetilde{X}$ and a subgroup $H$ of $G$ fixing $x$ we have a canonical split $H\to\widetilde{G}$ sending $h\in H$ to the unique pair $(\alpha\colon\tilde{X}\to\widetilde{X},h)$ such that $\alpha(\tilde{x})=\tilde{x}$.
@@
The $G$-equivariant Stolz exact sequence associated to $X$ is isomorphic to the $\widetilde{G}-$equivariant Stolz exact sequence associated to $\widetilde{X}$.
@@
For the paper at hand, the transition to the universal covering as in Proposition 3.6 is not really relevant. However, in other situations this point of view is really fruitful. To our knowledge, essentially all information about non-triviality of the groups $R^{spin}_{n}(X)$ and $Pos_{n}^{spin}(X)$ use higher index theory of the Dirac operator. A particularly transparent way to do this is to use the equivariant Dirac operator on the universal covering. This essentially means that one applies the isomorphism the groups “downstairs” for $X$ and the $\pi_{1}(X)$-equivariant groups “upstairs” for the universal covering $\tilde{X}$, as we have already mentioned in the introduction.
It is pretty obvious that the same observation does hold for non-free actions: even if we want to understand symmetric metrics of positive scalar curvature on a compact manifold $M$ with smooth action of a finite group $\Gamma$, it will be very useful to pass to the covering space which also sees the symmetries induced by the non-trivial fundamental group. This strategy has already been implemented in [XieYu, Section 5] and [XieYuZeidler]. The first paper constructs the map from the equivariant Stolz sequence to analysis (in form of the Higson-Roe exact sequence for the K-theory of the Roe $C^{*}$-algebra) in the sense of [PiazzaSchick]. The second paper then uses this to explicitly distinguish many bordism classes of equivariant positive scalar curvature metrics.
@@ 3.1 Refinements beyond bordism of positive scalar curvature metrics
Non-equivariantly, the situation here is very satisfactory: for a given connected closed spin manifold $M$ with $\dim(M)\geq 5$, if we use $X=B\pi_{1}M$ then $M$ itself admits a metric with positive scalar curvature if (and only if) the image of $u\colon M\to B\pi_{1}(M)]$ in $R_{n}^{spin}(B\pi_{1}(M))$ vanishes. Moreover, in this case $R_{n+1}^{spin}(B\pi_{1}(M))$ acts freely and transitively on the concordance classes of metrics of positive scalar curvature on $M$.
Of course, it would be very desirable to extend such results to the equivariant case. Unfortunately, the proof in the non-equivariant case uses as an important tool “handle cancellation” for Morse functions. It is known that equivariantly such cancellation is not even true. Therefore, a general treatment of the equivariant case seems not in reach at the moment. Under very special conditions on the action, positive existence results can be obtained. The strongest results in this direction we are aware of are obtained in [Hanke]. There, also the difficulties are discussed when one attempts to obtain more general results. In the paper at hand, we focus on the bordism context and do not attempt to contribute to obstruction and classification results for G-invariant metrics of positive scalar curvature on a fixed manifold $M$.
4 Invariance of R-groups under 2-equivalence
@@
Let $f\colon X\rightarrow Y$ be a continuous, 2-connected $G$-map between $G$-CW-complexes (in particular, it induces a bijection between components of fixed point sets and an isomorphism of the fundamental groups of all components of fixed point sets of $X$ and the corresponding ones of $Y$). Then for $\geq 6$ the functorially induced map
$f_{*}\colon R_{n}^{spin}(X)^{G}\rightarrow R_{n}^{spin}(Y)^{G}$
is an isomorphism.
@@
Surjectivity. We start by showing the surjectivity of the map
$f_{*}\colon R_{n}^{spin}(X)^{G}\rightarrow R_{n}^{spin}(Y)^{G}.$
Let us consider a class $W,\varphi\colon W\rightarrow Y,g]\in R_{n}^{spin}(Y)^{G}$. We want to find a bordant cycle whose reference map factors through $f$.
Consider W as a bordism between its boundary $\partial W$ and the empty set and choose a $G$-invariant Morse function $\alpha\colon W\to\mathbb{R}$ on it with critical points rearranged as described in [Milnor, Theorem 4.8], namely for any critical points $p_{i}$ and $p_{j}$ such that $f(p_{i})<f(p_{j})$, we have that $Ind_{f}(p_{i})<Ind_{f}(p_{j})$, where $Ind_{f}(p)$ denotes the Morse index at $p$ of the function $f$. Notice that we are going to use the enhanced version of this result to the equivariant setting, see for instance [Mayer] or [Wasserman].
Then there exists a suitable $t\in\mathbb{R}$ such that the subset $W_{1}:=\alpha^{-1}([0,t])\subset W$ consists only of $G$-handles of dimensions 0,1 and 2. We immediately obtain a decomposition of $W$ as $W_{1}\cup W_{2}$ such that $W_{1}$ is a bordism from the empty set to $M_{1}:=\alpha^{-1}(t)$ and $W_{2}$ a bordism from $M_{1}$ to $\partial W$. Of course, $W_{2}$ has only critical points $p_{i}$ with $Ind_{\alpha}(p_{i})\geq 3$. Consider now the function $\alpha$: this is a Morse function on $W_{2}$ seen as a bordism from $\partial W$ to $M_{1}$ with same critical points $p_{i}$ but with indices now given by $Ind_{-\alpha}(p_{i})=dim(W)-Ind_{\alpha}(p_{i})$. These critical points $p_{i}$ are then associated to $G$-equivariant $(Ind_{-\alpha}(p_{i})-1)$-surgeries, hence of codimension $Index_{\alpha}(p_{i})+1$ which is $\geq 3$.
This allows us to apply the Gromov-Lawson Theorem in its $G$-equivariant version as it is proved in [Hanke, Theorem 2]. This implies that we can extend the metric with positive scalar curvature $g$ on $\partial W$ to a $G$-invariant metric with positive scalar curvature $\bar{g}$ on $W_{2}$. Let us denote by $g_{1}$ its restriction to $M_{1}$. Observe that the triad $(W_{1},\varphi_{|W_{1}},g_{1})$ defines a class in $R_{n}^{spin}(Y)^{G}$ and the manifold $W\times[0,1]$ provides a bordism between $(W_{1},\varphi_{|W_{1}},g_{1})$ and $(W,\varphi,g)$.
@@
  • Since $f$ is 2-connected, up to $G$-homotopy we can assume that its restriction to the 2-skeleton
    $f^{(2)}\colon X^{(2)}\rightarrow Y^{(2)}\rightarrow Y$
    has a right inverse, i.e. there exists a $G$-equivariant map $h\colon Y^{(2)}\rightarrow X^{(2)}$ such that $f^{(2)}\circ h$ is homotopic to $i$. To see this, observe that the existence of such a map $h$ is guaranteed, up to $G$-homotopy, by Proposition 2.7. In fact, since $f^{(2)}$ is $2$-connected and $Y^{(2)}$ has dimension $\leq 2$, it suffices to apply 2.7 with $A=Y^{(2)}$, $B=X^{(2)}$, $C=Y$ and $f=f^{(2)}$ to the map $i\in[A,C]^{G}$. The surjectivity of $f_{*}$ then implies the existence of $h$.
@@
$f_{*}[W_{1},\psi\colon W_{1}\to X,g_{1}]=[W,\varphi\colon W\to Y,g]\in R^{spin}_{n}(Y)^{G},$
@@
Injectivity. In order to prove the injectivity of $f_{*}\colon R_{n}^{spin}(X)^{G}\to R_{n}^{spin}(Y)^{G}$, let us consider a class $W,\varphi:W\rightarrow X,g]\in R_{n}^{spin}(X)^{G}$ such that its image $f_{*}[W,\varphi:W\rightarrow X,g]$ is equal to the trivial element in $R_{n}^{spin}(Y)^{G}$. This means that there exists
  • an $n+1$-dimensional $G$-manifold with corners $B$, whose codimension 1 faces are $W$ itself together with a bordism $V$ from $\partial W$ to the empty set, which intersect in the only codimension 2 face $\partial W=W\cap V$;
@@
By construction, the bordism $B_{2}$ is the trace of surgeries of codimension $\geq 3$. Therefore, we can apply the equivariant version [Hanke, Theorem 2] of the Gromov-Lawson Theorem to extend the metric $g_{\bar{V}}$ to a $G$-invariant metric of positive scalar curvature $g_{2}$ on $B_{2}$. Let us denote by $g_{1}$ the $G$ -invariant metric of positive scalar curvature obtained by restricting $g_{2}$ to $W_{1}$.
The last fact to prove is that $\Psi_{|B_{1}}\colon B_{1}\to Y$ factors through $f\colon X\to Y$. Indeed $B_{1}$ is obtained form $W$ by attaching, up to homotopy, cells of dimension up to 2. Define a map $h\colon B_{1}\times\{0\}\cup W\times[0,1]\to Y$ as the restriction of $\Psi|_{B_{1}}\circ pr$ to this subspace of $B_{1}\times[0,1]$, where $pr\colon B_{1}\times[0,1]\to B_{1}$ is the obvious projection. Note that by assumption the restriction of this map to $W\times\{1\}$ equals $f\circ\varphi$. These are exactly the conditions of the relative precursor [ttd, (2.5) Proposition] of Proposition 2.7 which now implies in particular (because $f$ is $2$-connected) that there exists an extension $K\colon B_{1}\to X$ of $\varphi$.
Now observe that $(B_{1},\Phi,g_{1})$ is a bordism between $(W,\varphi,g)$ and the trivial cycle $(W_{1},K|_{W_{1}},g_{1}|_{\partial W})$ in $R^{spin}_{n}(X)^{G}$ (trivial because the metric extends to the metric $g_{1}$ of positive scalar curvature on all of $W_{1}$), the injectivity of $f_{*}$ is proved.
@@ 5 Universal Spaces
It is a known fact that $B\Gamma$ is characterized by the fact that it is connected, has fundamental group $\Gamma$, whereas all higher homotopy groups are trivial.
@@
Define for a family of subgroup $\mathcal{F}$ of $G$ the following orbit category
$\operatorname{Orb}(G,\mathcal{F})$
whose objects are all subgroups in $\mathcal{F}$ and morphisms from $H\in\mathcal{F}$ to $K\in\mathcal{F}$ are $G$-maps $G/H\to G/K$ for $H,K\in\mathcal{F}$. Note [^2] To be explicit: as $G/H$ consists of a single $G$-orbit, the map is determined by the image of the coset $H$. Say this image is the coset $gK$. Then by $G$-equivariance for each $x\in G$ we must have that $xH$ is mapped to $xgK$. To be well defined, for every $x\in G$ and $h\in H$ we require that $xhgK$ and $xgK$ are in the same $K$-coset, i.e. $g^{-1}x^{-2}xhg\in K$ for each $h\in H$, which is the condition $g^{-1}Hg\subset K$. that any such map is of the form $xH\mapsto xgK$ for a well defined coset $g]\in G/K$, where we also have that $g^{-1}Hg\subset K$ —precisely the condition for this map to be well defined.
@@
$\Pi_{1}(X;G)\colon\operatorname{Orb}(G,\mathcal{I}(X))\to Groupoids$
@@
  • to a group $H\in\mathcal{I}(X)$ (the isotropy family) the fundamental groupoid of $X^{H}$ restricted to the 0-skeleton of $X^{H}$ (meaning that we take the full subgroupoid), which we denote by $\Pi_{1}(X^{H})_{|X^{H}_{(0)}}$,
  • to a morphism from $H$ to $K$ in the category $\operatorname{Orb}(G,\mathcal{I}(X))$ given as $G$-map $G/H\mapsto G/K;xH\mapsto xgK$ the morphism of groupoids between $\Pi_{1}(X^{K})_{|X^{K}_{(0)}}$ and $\Pi_{1}(X^{H})_{|X^{H}_{(0)}}$ induced by the map $X^{K}\to X^{H}$ defined as $x\mapsto gx$.
@@
$\Pi_{1}(X;G)\colon\operatorname{Orb}(G)\to Groupoids,$
assigning to $H\notin\mathcal{I}(X)$ the empty groupoid. Note that the empty groupoid has a unique map to any other groupoid, but is not the target of any map from a non-empty groupoid. The former is good because it determines the value of $\Pi_{1}(X;G)$ on morphisms in $\operatorname{Orb}(G)$ with domain $H$. The latter is no problem because families are closed under conjugation and under taking subgroups and therefore if $K\in\mathcal{I}(X)$ and $H\notin\mathcal{I}(X)$ there never is a morphism from $K$ to $H$ in $\mathcal{ALL}$333recall that a morphism from $H$ to $K$ exists if and only if $\exists g\in G$ with $g^{-1}Hg\subset K$.
@@
More conceptually, to construct the fundamental groupoid functor, we can observe that we have the canonical morphism
$\mathrm{map}^{G}(G/H,X)\to X^{H}$
between the mapping space of $G$-equivariant maps and the fixed point set, sending $f\colon G/H\to X$ to $f(eH)$. The functoriality is then just given by precomposition.
Observe that the construction of the fundamental groupoid functor is itself functorial. This means that if $\varphi\colon Y\to X$ is a $G$-equivariant cellular map between $G$-CW-complexes, then there is an induced natural transformation
$\varphi_{\#}\colon\Pi_{1}(Y;G)\to\Pi_{1}(X;G)$
whose component at $H$ is the homomorphism of groupoids
$\varphi_{\#}(H)\colon\Pi_{1}(Y^{H})_{|Y^{H}_{(0)}}\to\Pi_{1}(X^{H})_{|X^{H}_{(0)}}$
induced by $\varphi_{|Y^{H}}\colon Y^{H}\to X^{H}$.

### v3

@@ 3 The Stolz exact sequence
  1. (1)
    For every component $X_{c}$ of $X$ consider the induced $G$-invariant subspace $G\cdot X_{c}$. We then have a disjoint union decomposition into $G$-invariant subsets $X=\bigsqcup G\cdot X_{c}$ with $X_{c}$ connected (choosing one representative in each $G$-orbit of components). Clearly, every group in the Stolz exact sequence and the whole sequence then split canonically as a direct product, with one factor for each $G\cdot X_{c}$.
  2. (2)
    For a $G$-space $Y=G\cdot X_{c}$ with $X_{c}$ connected, consider the subgroup $G_{0}\subset G$ of all elements which map $X_{c}$ to itself. Then $X_{c}$ is a $G_{0}$-space and $Y$ is obtained by “induction”: $Y=G\times_{G_{0}}X_{c}$. Whenever we have a $G$-map $f\colon M\to Y$ which could be part of a cycle or a bordism for the groups in the Stolz exact sequence, then we obtain $M_{0}:=f^{-1}(X_{c})$} a union of components of $M$ on which $G_{0}$ acts (restricting the action of $G$ on $M$). Then $f|_{M_{0}}\colon M_{0}\to X_{c}$} is a $G_{0}$-equivariant map and we obtain $M$ and $f$ by induction: $M=G\times_{G_{0}}M_{0}$ and $f=\operatorname{id}_{G}\times_{G_{0}}f|_{M_{0}}$. It follows that induction gives an isomorphism (already on the level of cycles and relations)
    $\mathrm{ind}_{G_{0}}^{G}\colon\ R^{\rm spin}_{*}(X_{c})^{G_{0}}\to R^{\rm spin}_{*}(Y)^{G}$
    and the same for the other groups in the sequence of Proposition 3.3 and the maps between them.
@@
This again follows from covering theory: if we fix $x\in X$ with lift $\tilde{x}\in\tilde{X}$ and a subgroup $H$ of $G$ fixing $x$ we have a canonical split $H\to\widetilde{G}$ sending $h\in H$ to the unique pair $\bigl(\alpha\colon\tilde{X}\to\tilde{X},h\bigr)$ such that $\alpha(\tilde{x})=\tilde{x}$.
@@
The $G$-equivariant Stolz exact sequence associated to $X$ is isomorphic to the $\widetilde{G}-$equivariant Stolz exact sequence associated to $\tilde{X}$.
@@
For the paper at hand, the transition to the universal covering as in Proposition 3.6 is not really relevant. However, in other situations this point of view is really fruitful. To our knowledge, essentially all information about non-triviality of the groups $R^{\rm spin}_{n}(X)$ and $\mathrm{Pos}_{n}^{\rm spin}(X)$ use higher index theory of the Dirac operator. A particularly transparent way to do this is to use the equivariant Dirac operator on the universal covering. This essentially means that one applies the isomorphism between the groups “downstairs” for $X$ and the $\pi_{1}(X)$-equivariant groups “upstairs” for the universal covering $\tilde{X}$, as we have already mentioned in the introduction.
It is pretty obvious that the same observation does hold for non-free actions: even if we want to understand symmetric metrics of positive scalar curvature on a compact manifold $M$ with smooth action of a finite group $\Gamma$, it will be very useful to pass to the covering space which also sees the symmetries induced by the non-trivial fundamental group. This strategy has already been implemented in [14, Section 5] and [15]. The first paper constructs the map from the equivariant Stolz sequence to analysis (in form of the Higson–Roe exact sequence for the K-theory of the Roe $C^{*}$-algebra) in the sense of [5]. The second paper then uses this to explicitly distinguish many bordism classes of equivariant positive scalar curvature metrics.
@@ 3.1 Refinements beyond bordism of positive scalar curvature metrics
Non-equivariantly, the situation here is very satisfactory: for a given connected closed spin manifold $M$ with $\dim(M)\geq 5$, if we use $X=B\pi_{1}M$ then $M$ itself admits a metric with positive scalar curvature if (and only if) the image of $u\colon M\to B\pi_{1}(M)]$ in $R_{n}^{\rm spin}(B\pi_{1}(M))$ vanishes. Moreover, in this case $R_{n+1}^{\rm spin}(B\pi_{1}(M))$ acts freely and transitively on the concordance classes of metrics of positive scalar curvature on $M$.
Of course, it would be very desirable to extend such results to the equivariant case. Unfortunately, the proof in the non-equivariant case uses as an important tool “handle cancellation” for Morse functions. It is known that equivariantly such cancellation is not even true. Therefore, a general treatment of the equivariant case seems not in reach at the moment. Under very special conditions on the action, positive existence results can be obtained. The strongest results in this direction we are aware of are obtained in [2]. There, also the difficulties are discussed when one attempts to obtain more general results. In the paper at hand, we focus on the bordism context and do not attempt to contribute to obstruction and classification results for G-invariant metrics of positive scalar curvature on a fixed manifold $M$.
4 Invariance of $\boldsymbol{R}$-groups under 2-equivalence
@@
Let $f\colon X\rightarrow Y$ be a continuous, $2$-connected $G$-map between $G$-CW-complexes ((in particular, it induces a bijection between components of fixed point sets and an isomorphism of the fundamental groups of all components of fixed point sets of $X$ and the corresponding ones of $Y)$. Then for $\geq 6$ the functorially induced map $f_{*}\colon R_{n}^{\rm spin}(X)^{G}\rightarrow R_{n}^{\rm spin}(Y)^{G}$ is an isomorphism.
@@
Surjectivity. We start by showing the surjectivity of the map $f_{*}\colon R_{n}^{\rm spin}(X)^{G}\!\rightarrow\!R_{n}^{\rm spin}(Y)^{G}$. Let us consider a class $W,\varphi\colon W\rightarrow Y,g]\in R_{n}^{\rm spin}(Y)^{G}$. We want to find a bordant cycle whose reference map factors through $f$.
Consider W as a bordism between its boundary $\partial W$ and the empty set and choose a $G$-invariant Morse function $\alpha\colon W\to\mathbb{R}$ on it with critical points rearranged as described in [4, Theorem 4.8], namely for any critical points $p_{i}$ and $p_{j}$ such that $f(p_{i})<f(p_{j})$, we have that $\operatorname{Ind}_{f}(p_{i})<\operatorname{Ind}_{f}(p_{j})$}, where $\operatorname{Ind}_{f}(p)$ denotes the Morse index at $p$ of the function $f$. Notice that we are going to use the enhanced version of this result to the equivariant setting, see for instance [3] or [13].
Then there exists a suitable $t\in\mathbb{R}$ such that the subset $W_{1}:=\alpha^{-1}([0,t])\subset W$ consists only of $G$-handles of dimensions 0, 1 and 2. We immediately obtain a decomposition of $W$ as $W_{1}\cup W_{2}$} such that $W_{1}$ is a bordism from the empty set to $M_{1}:=\alpha^{-1}(t)$ and $W_{2}$ a bordism from $M_{1}$ to $\partial W$. Of course, $W_{2}$ has only critical points $p_{i}$ with $\operatorname{Ind}_{\alpha}(p_{i})\geq 3$. Consider now the function $\alpha$: this is a Morse function on $W_{2}$ seen as a bordism from $\partial W$ to $M_{1}$ with same critical points $p_{i}$ but with indices now given by $\operatorname{Ind}_{-\alpha}(p_{i})=\dim(W)-\operatorname{Ind}_{\alpha}(p_{i})$. These critical points $p_{i}$ are then associated to $G$-equivariant $(\operatorname{Ind}_{-\alpha}(p_{i})-1)$}-surgeries, hence of codimension $\operatorname{Index}_{\alpha}(p_{i})+1$} which is $\geq 3$.
This allows us to apply the Gromov–Lawson theorem in its $G$-equivariant version as it is proved in [2, Theorem 2]. This implies that we can extend the metric with positive scalar curvature $g$ on $\partial W$ to a $G$-invariant metric with positive scalar curvature $\bar{g}$ on $W_{2}$. Let us denote by $g_{1}$ its restriction to $M_{1}$. Observe that the triad $(W_{1},\varphi_{|W_{1}},g_{1})$ defines a class in $R_{n}^{\rm spin}(Y)^{G}$ and the manifold $W\times[0,1]$ provides a bordism between $(W_{1},\varphi_{|W_{1}},g_{1})$ and $(W,\varphi,g)$.
@@
  • Since $f$ is 2-connected, up to $G$-homotopy we can assume that its restriction to the 2-skeleton $f^{(2)}\colon\ X^{(2)}\rightarrow Y^{(2)}\rightarrow Y$ has a right inverse, i.e., there exists a $G$-equivariant map $h\colon Y^{(2)}\rightarrow X^{(2)}$} such that $f^{(2)}\circ h$ is homotopic to $i$. To see this, observe that the existence of such a map $h$ is guaranteed, up to $G$-homotopy, by Proposition 2.7. In fact, since $f^{(2)}$ is $2$-connected and $Y^{(2)}$ has dimension $\leq 2$, it suffices to apply Proposition 2.7 with $A=Y^{(2)}$, $B=X^{(2)}$, $C=Y$ and $f=f^{(2)}$ to the map $i\in[A,C]^{G}$. The surjectivity of $f_{*}$ then implies the existence of $h$.
@@
$f_{*}[W_{1},\psi\colon W_{1}\to X,g_{1}]=[W,\varphi\colon W\to Y,g]\in R^{\rm spin}_{n}(Y)^{G},$
@@
Injectivity. In order to prove the injectivity of $f_{*}\colon R_{n}^{\rm spin}(X)^{G}\to R_{n}^{\rm spin}(Y)^{G}$, let us consider a class $W,\varphi\colon W\rightarrow X,g]\in R_{n}^{\rm spin}(X)^{G}$ such that its image $f_{*}[W,\varphi\colon W\rightarrow X,g]$ is equal to the trivial element in $R_{n}^{\rm spin}(Y)^{G}$. This means that there exists
  • an $(n+1)$-dimensional $G$-manifold with corners $B$, whose codimension 1 faces are $W$ itself together with a bordism $V$ from $\partial W$ to the empty set, which intersect in the only codimension 2 face $\partial W=W\cap V$;
@@
By construction, the bordism $B_{2}$ is the trace of surgeries of codimension $\geq 3$. Therefore, we can apply the equivariant version [2, Theorem 2] of the Gromov–Lawson theorem to extend the metric $g_{\bar{V}}$ to a $G$-invariant metric of positive scalar curvature $g_{2}$ on $B_{2}$. Let us denote by $g_{1}$ the $G$ -invariant metric of positive scalar curvature obtained by restricting $g_{2}$ to $W_{1}$.
The last fact to prove is that $\Psi_{|B_{1}}\colon B_{1}\to Y$ factors through $f\colon X\to Y$. Indeed, $B_{1}$ is obtained form $W$ by attaching, up to homotopy, cells of dimension up to 2. Define a map $h\colon B_{1}\times\{0\}\cup W\times[0,1]\to Y$ as the restriction of $\Psi|_{B_{1}}\circ\mathrm{pr}$ to this subspace of $B_{1}\times[0,1]$, where $\mathrm{pr}\colon B_{1}\times[0,1]\to B_{1}$} is the obvious projection. Note that by assumption the restriction of this map to $W\times\{1\}$ equals $f\circ\varphi$. These are exactly the conditions of the relative precursor [12, Proposition 2.5] of Proposition 2.7 which now implies in particular (because $f$ is $2$-connected) that there exists an extension $K\colon B_{1}\to X$ of $\varphi$.
Now observe that $(B_{1},\Phi,g_{1})$ is a bordism between $(W,\varphi,g)$ and the trivial cycle $(W_{1},K|_{W_{1}},\allowbreak g_{1}|_{\partial W})$ in $R^{\rm spin}_{n}(X)^{G}$ (trivial because the metric extends to the metric $g_{1}$ of positive scalar curvature on all of $W_{1}$), the injectivity of $f_{*}$ is proved. ∎
@@ 5 Universal spaces
It is known that $B\Gamma$ is characterized by the fact that it is connected, has fundamental group $\Gamma$, whereas all higher homotopy groups are trivial.
@@
Define for a family of subgroup $\mathcal{F}$ of $G$ the following orbit category $\operatorname{Orb}(G,\mathcal{F})$, whose objects are all subgroups in $\mathcal{F}$ and morphisms from $H\in\mathcal{F}$ to $K\in\mathcal{F}$ are $G$-maps $G/H\to G/K$ for $H,K\in\mathcal{F}$. Note [^2] To be explicit: as $G/H$ consists of a single $G$-orbit, the map is determined by the image of the coset $H$. Say this image is the coset $gK$. Then by $G$-equivariance for each $x\in G$ we must have that $xH$ is mapped to $xgK$. To be well defined, for every $x\in G$ and $h\in H$ we require that $xhgK$ and $xgK$ are in the same $K$-coset, i.e., $g^{-1}x^{-2}xhg\in K$ for each $h\in H$, which is the condition $g^{-1}Hg\subset K$. that any such map is of the form $xH\mapsto xgK$ for a well defined coset $g]\in G/K$, where we also have that $g^{-1}Hg\subset K$ —precisely the condition for this map to be well defined.
@@
$\Pi_{1}(X;G)\colon\ \operatorname{Orb}(G,\mathcal{I}(X))\to\text{groupoids},$
@@
  • to a group $H\in\mathcal{I}(X)$ (the isotropy family) the fundamental groupoid of $X^{H}$ restricted to the 0-skeleton of $X^{H}$ (meaning that we take the full subgroupoid), which we denote by $\Pi_{1}\bigl(X^{H}\bigr)_{|X^{H}_{(0)}}$,
  • to a morphism from $H$ to $K$ in the category $\operatorname{Orb}(G,\mathcal{I}(X))$ given as $G$-map $G/H\mapsto G/K;xH\mapsto xgK$ the morphism of groupoids between $\Pi_{1}(X^{K})_{|X^{K}_{(0)}}$ and $\Pi_{1}\bigl(X^{H}\bigr)_{|X^{H}_{(0)}}$ induced by the map $X^{K}\to X^{H}$ defined as $x\mapsto gx$.
@@
$\Pi_{1}(X;G)\colon\ \operatorname{Orb}(G)\to\text{groupoids},$
assigning to $H\notin\mathcal{I}(X)$ the empty groupoid. Note that the empty groupoid has a unique map to any other groupoid, but is not the target of any map from a non-empty groupoid. The former is good because it determines the value of $\Pi_{1}(X;G)$ on morphisms in $\operatorname{Orb}(G)$ with domain $H$. The latter is no problem because families are closed under conjugation and under taking subgroups and therefore if $K\in\mathcal{I}(X)$ and $H\notin\mathcal{I}(X)$ there never is a morphism from $K$ to $H$ in $\mathcal{ALL}$. [^3] Recall that a morphism from $H$ to $K$ exists if and only if $\exists g\in G$ with $g^{-1}Hg\subset K$.
@@
More conceptually, to construct the fundamental groupoid functor, we can observe that we have the canonical morphism $\mathrm{map}^{G}(G/H,X)\to X^{H}$ between the mapping space of $G$-equivariant maps and the fixed point set, sending $f\colon G/H\to X$ to $f(eH)$. The functoriality is then just given by precomposition.
Observe that the construction of the fundamental groupoid functor is itself functorial. This means that if $\varphi\colon Y\to X$ is a $G$-equivariant cellular map between $G$-CW-complexes, then there is an induced natural transformation $\varphi_{\#}\colon\Pi_{1}(Y;G)\to\Pi_{1}(X;G)$ whose component at $H$ is the homomorphism of groupoids $\varphi_{\#}(H)\colon\Pi_{1}\bigl(Y^{H}\bigr)_{|Y^{H}_{(0)}}\to\Pi_{1}\bigl(X^{H}\bigr)_{|X^{H}_{(0)}}$ induced by $\varphi_{|Y^{H}}\colon Y^{H}\to X^{H}$}.